# A Practical Adaptive Subgame Perfect Gradient Method

### v1

@@ 1 Introduction
There is a common trade-off in optimization between strong convergence guarantees and practical speed. For first-order convex optimization, linesearch methods, in particular quasi-Newton methods, are well-established and popular algorithms. When measured by practical performance, L-BFGS [1] and its variants (e.g., [2, 3]) are state-of-the-art for many smooth convex problems, often converging superlinearly. However, despite this empirical performance, these methods have rather limited non-asymptotic guarantees. At the other extreme, accelerated gradient methods with comprehensive convex optimization theory are often outperformed by these linesearch methods. For example, the Optimized Gradient Method (OGM) [4] is known to provide the minimax optimal convergence guarantee for smooth convex problems but can fail to converge faster than its worst-case rate even on very simple instances [^3] See the example of minimizing $f(x)=\tfrac{1}{2}x^{2}$ in [5] where OGM attains the minimax worst-case rate..
@@ 1.1 Related Literature
where $\mathcal{F}_{L}$ denotes the family of all $L$-smooth convex problems with a minimizer $x_{\star}$ and initialization $x_{0}$. Note OGM is not adaptive or parameter-free as it relies on knowledge of the smoothness constant $L$. OGM and its theory are formally introduced in Section 2.1.1.
@@
Quasi-Newton methods [30, 31, 32, 33, 1, 2, 3] attempt to find a local minimum for a function by applying Newton’s method with an approximation of the inverse Hessian. At each iteration $n$, one performs a low-rank update to the inverse Hessian approximation, $B_{n}$, and a search direction is computed as $v=-B_{n}\nabla f(x_{n})$. Once a descent direction is found, a linesearch method is applied to determine a stepsize and update the current iterate. Common options for linesearch methods include Armijo backtracking [34, 35], Strong-Wolfe conditions [36, 35], and a modified robust method introduced by Hager and Zhang in [37]. In L-BFGS [1], instead of storing the full matrix $B_{n}$, one stores only a limited memory of past data. This allows the matrix update and search direction calculation to be performed using only vector-vector products, preventing any large matrix storage. While L-BFGS performs extremely well in practice and often has superlinear convergence in a neighborhood of the solution, it only provides limited non-asymptotic guarantees on solution quality.
@@
Preconditioning can be viewed as selecting an inner product via a preconditioner $B$ to reduce the problem’s difficulty, with respect to some measure. In smooth convex problems, a good preconditioner would reduce $\frac{L_{B}}{2}\|x_{0}-x_{\star}\|_{B}^{2}$ as this controls convergence (where $L_{B}$ is a global smoothness constant of $f$ with respect to a modified inner product $\langle\cdot,\cdot\rangle_{B}$). In smooth strongly convex problems, one would like a preconditioner $B$ minimizing the condition number $L_{B}/\mu_{B}$. For $f\in C^{2}$, selecting $B$ as the inverse of $\nabla^{2}f(x_{\star})$ drives this condition number to one in a neighborhood of $x_{\star}$. One can view quasi-Newton methods then as gradient descent with its preconditioner being adjusted at runtime to roughly approximate this inverse Hessian. The work [20] introduces a multidimensional backtracking technique to update a diagonal preconditioner at runtime. As a modern advancement of this idea, the Online Gradient Scaling Method of [19] uses the online learning regret bound toolbox to adjust the preconditioner at runtime. Asymptotically, this ensures an optimal preconditioned performance measure for gradient descent.
@@
While effective, a downside of backtracking is that it can be “wasteful” with respect to oracle evaluations. When a backtrack occurs, the new oracle information computed at the candidate iterate is effectively discarded. Linesearch-free methods manage to avoid this small inefficiency through various techniques. In bundle methods [21, 22, 23, 24], this is handled with a “null step” in which the function value and gradient information are still saved in the bundle to inform future iterations. Starting with the introduction of adaptive gradient descent by [9], many recent works [8, 10, 11, 13, 16, 17, 18] adaptively learn the smoothness constant without backtracking, while also providing a non-asymptotic convergence guarantee. The task of designing methods that estimate an unknown level of strong convexity is typically much harder than adapting to smoothness. Recently, [12] introduced a method for adaptively learning strong convexity constants while guaranteeing (suboptimal) linear global convergence. Our ASPGM will employ a similar estimate as a heuristic to benefit from strong convexity.
@@ 1.2 Our Contributions
  • An Adaptive Subgame Perfect Gradient Method (ASPGM). The strong non-asymptotic convergence guarantees and certificates resulting from BSPGM’s subgame perfection provide a strong restarting condition for the algorithm, ensuring a contraction under strong convexity. Combining this restarting with preconditioning gives our full algorithm, ASPGM.
@@ 2 Preliminaries
where $f$ is convex and differentiable with gradient $\nabla f$ locally Lipschitz and with a minimizer $x_{\star}$. We highlight the fact that a global Lipschitz constant on the gradient $L$ need not be known, or even exist, for our algorithm to be well-defined and progress. We denote $I$ as the identity matrix, $e_{i}$ as the $i$th standard basis vector, and $\mathbf{1}$ as the all ones vector, where the dimension of each should be clear from context.
@@
$\tilde{L}_{B}(x,y)=\frac{f(y)-f(x)-\langle{\nabla_{\!\!{\scriptscriptstyle B}}f(x)},{y-x}\rangle_{B}}{\frac{1}{2}\|{\nabla_{\!\!{\scriptscriptstyle B}}f(x)-\nabla_{\!\!{\scriptscriptstyle B}}f(y)}\|_{B}^{2}}$ (2.5)
@@ 2.1.1  OGM - An Inductive View
for some $\tau_{n}>0$ and iterate sequences $x_{n},z_{n+1}$. The base case of this induction can be established for any given $x_{0}$ by setting $z_{1}=x_{0}-\frac{2}{L}g_{0},\ \tau_{0}=2$ since we then have $H_{0}=Q_{\star,0}(L)\geq 0.$ Then OGM iterates by setting
@@ 2.1.2  SPGM - A Dynamically Optimized Induction
$q_{i}$ $:=\tau_{i}(f_{i}-\frac{1}{2L}\|{g_{i}}\|_{B}^{2})+\frac{L}{2}\|{z_{i+1}}\|_{B}^{2}-\frac{L}{2}\|{x_{0}}\|_{B}^{2}-v_{m}\tau_{i}-\langle{L(z_{i+1}-x_{0})},{x_{0}}\rangle_{B}$
$r_{i}$ $:=f_{i}-\langle{g_{i}},{x_{i}-x_{0}}\rangle_{B}-\frac{1}{2L}\|{g_{i}}\|_{B}^{2}-v_{m}.$
@@ 2.1.3  OBL - A Backtracking-Friendly Induction
Both of the above methods relied on knowledge of the global smoothness constant $L$, which can be both unrealistic in practice and lead to overly conservative algorithms. One classic approach to avoiding this is the use of backtracking, where at each iteration an estimate $L_{n}$ is utilized to compute a step. Then, whichever inequalities are needed by the proof are checked; if they hold the step is accepted, otherwise $L_{n}$ is increased, typically to $2L_{n}$, and the process repeats.
@@
Following this, one has that (see [38, Theorem 6])
@@ 3.1 Dynamically Optimized Induction
We begin by introducing some notation. Given a vector $\tau=(\tau_{n-k},\dots,\tau_{n-1})$ and $L_{n}>0$, define the vector $v\in\mathbb{R}^{k}$ by its components:
@@
for $i\in[n-k,n-1]$. Next, letting $J=\{i\in[n-k,n-1]\mid\tau_{i}>0\}$, define two special indices $m$ and $s$ according to
@@
$a_{i}$ $:=\tau_{i}(f_{i}-\frac{1}{2L_{i}}\|{g_{i}}\|_{B}^{2})+\frac{L_{i}}{2}\|{z_{i+1}}\|_{B}^{2}-\frac{L_{i}}{2}\|{x_{0}}\|_{B}^{2}-v_{m}\tau_{i}-\langle{L_{i}(z_{i+1}-x_{0})},{x_{0}}\rangle_{B}$ (3.5)
@@
Let $Z\in\mathbb{R}^{d\times k}$ be the matrix with columns now given by $\frac{L_{i}}{L_{n}}(z_{i+1}-x_{0})$ and indexed by [nk,n1][n-k,n-1]. Similarly, let $G\in\mathbb{R}^{d\times k}$ be the matrix with columns now given by $\frac{g_{i}}{L_{n}}$ and indexed by [nk,n1][n-k,n-1]. Once again, by rearranging (3.8) and applying a bound on $\Delta^{\prime}$, we find that the following constraints must be satisfied:
@@
Observe that maximizing $\tau^{\prime}$ in (3.12) once again amounts to solving a low-dimensional second order cone program (SOCP). We emphasize that this SOCP is independent of the problem dimension $d$. Specifically, the subproblem (3.12) has dimension $2k$, where $k$ is the chosen memory size. By appropriately tracking and storing $O(k^{2})$ past vector-vector products (i.e., $Z^{T}Z,G^{T}G,G^{T}Z$), we eliminate any matrix-vector products with full-dimension matrices. Additionally, the problem has a linear objective with a single quadratic constraint, further reducing the complexity of the subproblem. Thus, for high-dimensional ($k\ll d$) problems with moderate costs for gradient and/or function evaluations, the per iteration computational cost of (3.12) becomes negligible.
@@ 3.2 A Backtracking-free Subgame Perfect Gradient Method (BSPGM)
This dynamic induction already gives way to a reasonable backtracking method. Given $\tau^{\prime}$ and $z^{\prime}$ after solving (3.12), generate $x_{n}$ (along with $\tau_{n},z_{n+1},\Delta_{n}$) according to (2.16). We then check if the current estimate of $L_{n}$ is valid, i.e., if $Q_{m,n}(L_{n})\geq 0$. If so, accept $x_{n}$ and continue. If not, reject $x_{n}$, increment $L_{n}\leftarrow 2L_{n}$, and repeat the update rule. Note that when $x_{n}$ is accepted, this indicates that our hypothesis $U_{n}\geq 0$ is valid, so the induction can continue.
While this approach represents an effective algorithm, it is inefficient in terms of how it uses new oracle information. When a proposed next iterate $x_{n}$ is rejected, the new gradient $g_{n}$ and function value $f_{n}$ are effectively discarded without providing additional information other than the fact that the current estimate $L_{n}$ is too small. For high-dimensional problems, first-order oracles can have a large computational cost; we therefore want to learn as much information as possible from each oracle call.
@@
This strategy is similar to the typical bundle method approach of allowing both “serious steps” where the main iterate sequence updates and “null steps” where only the bundle of first-order information is improved. Algorithm˜4 formalizes our use of this conditional approach. We emphasize that while there are remnants of a backtracking mechanism in how we iteratively increase $L_{n}$, this method is notionally distinct as we are not discarding any computed $x_{n}$ and associated first-order oracle values when $L_{n}$ increases. Rather, $(x_{n},f_{n},g_{n})$ are incorporated into our bundle of historical information via the nonnegative quantity $W_{\star,n}$. We also note the use of $\tilde{L}_{B}(x_{m},x_{n})$ to increment $L_{n}$ matches the auto-conditioning approach used by [9, 11, 8] and others.
@@ 3.3 Convergence Guarantees and Subgame Perfection of BSPGM
To maintain our inductive hypothesis $U_{n}\geq 0$, we simply have to verify that
@@
$f(x_{n})-\frac{1}{2L_{n}}\|g_{n}\|^{2}_{B}-f(x_{\star})$ $\leq\frac{L_{n}\|x_{0}-x_{\star}\|^{2}_{B}+\Delta_{n}}{2\tau_{n}}\qquad\qquad$ $\text{if\ }n<N$ (3.14)
$f(x_{N})-f(x_{\star})$ $\leq\frac{L_{N}\|x_{0}-x_{\star}\|^{2}_{B}+\Delta_{N}}{2\tau_{N}}\qquad\qquad$ $\text{if\ }n=N.$ (3.15)
@@
The final bound above is minimax optimal among all gradient-span methods for minimization against such an oracle and is also attained by the simpler OBL method, see [38, Theorem 4]. That is, letting $\mathcal{A}$ denote the set of gradient span methods producing a final iterate after $N$ iterations,
@@ 3.4 Adaptive Restart
${\tau}^{(\ell)}_{n}\geq\frac{2{L}^{(\ell)}_{n}}{{\mu}^{(\ell)}}+\frac{{L}^{(\ell)}_{n}{\Delta}^{(\ell)}_{n}}{f({x}^{(\ell)}_{0})-f({x}^{(\ell)}_{n})},\qquad\qquad f({x}^{(\ell)}_{0})-f({x}^{(\ell)}_{n})>0.$ (3.16)
@@
$n\geq\sqrt{\frac{4{L}^{(\ell)}_{n}}{\mu}+\frac{2{L}^{(\ell)}_{n}{\Delta}^{(\ell)}_{n}}{f({x}^{(\ell)}_{0})-f(x_{\star})}}+\log_{2}\left(\frac{{L}^{(\ell)}_{n}}{{L_{0}}^{(\ell)}}\right).$
@@
In practice, accurately estimating ${\mu}^{(\ell)}$ (at each epoch) can be difficult and lead algorithms to be overly conservative. Numerically, we found strong practical performance by constructing a live estimate of $\mu$ similar to the approach of [12]. We initially set ${\mu}^{(\ell)}_{0}$$\infty$ and update at each iteration according to

### v2

@@ 1 Introduction
There is a common trade-off in optimization between strong convergence guarantees and practical speed. For first-order convex optimization, linesearch methods, in particular quasi-Newton methods, are well-established and popular algorithms. When measured by practical performance, L-BFGS [43] and its variants (e.g., [54, 48]) are state-of-the-art for many smooth convex problems, often converging superlinearly. However, despite this empirical performance, these methods have rather limited non-asymptotic guarantees. At the other extreme, accelerated gradient methods with comprehensive convex optimization theory are often outperformed by these linesearch methods. For example, the Optimized Gradient Method (OGM) [23] is known to provide the minimax optimal convergence guarantee for smooth convex problems but can fail to converge faster than its worst-case rate even on very simple instances [^3] See the example of minimizing $f(x)=\tfrac{1}{2}x^{2}$ in [19] where OGM fails to accelerate beyond the minimax worst-case rate..
@@ 1.1 Related Literature
where $\mathcal{F}_{L}$ denotes the family of all $L$-smooth convex problems with a minimizer $x_{\star}$ and initialization $x_{0}$. Note that OGM is not adaptive or parameter-free as it relies on knowledge of the smoothness constant $L$. OGM and its theory are formally introduced in Section 2.1.1.
@@
Quasi-Newton methods [6, 15, 18, 49, 43, 54, 48] attempt to find a minimizer for a function by applying Newton’s method with an approximation of the inverse Hessian. At each iteration $n$, one performs a low-rank update to the inverse Hessian approximation, $B_{n}$, and a search direction is computed as $v=-B_{n}\nabla f(x_{n})$. Once a descent direction is found, a linesearch method is applied to determine a stepsize and update the current iterate. Common options for linesearch methods include Armijo backtracking [2, 42], Strong-Wolfe conditions [52, 42], and a modified robust method introduced by Hager and Zhang in [20]. In L-BFGS [43], instead of storing the full matrix $B_{n}$, one stores only a limited memory of past data. This allows the matrix update and search direction calculation to be performed using only vector-vector products, avoiding the storage of dense matrices. While L-BFGS performs extremely well in practice and often has (super)linear convergence in a neighborhood of the solution, it only provides limited non-asymptotic guarantees on solution quality [^4] Although recent progress [21, 47, 22] has been made establishing non-asymptotic guarantees for standard BFGS with additional assumptions on the objective $f$..
@@
Preconditioning can be viewed as selecting an inner product via a preconditioner $B$ to reduce the problem’s difficulty, with respect to some measure. In smooth convex problems, a good preconditioner would reduce $\frac{L_{B}}{2}\|x_{0}-x_{\star}\|_{B}^{2}$ as this controls convergence (where $L_{B}$ is a global smoothness constant of $f$ with respect to a modified inner product $\langle\cdot,\cdot\rangle_{B}$). In smooth strongly convex problems, one would like a preconditioner $B$ minimizing the condition number $L_{B}/\mu_{B}$. For $f\in C^{2}$, selecting $B$ as the inverse of $\nabla^{2}f(x_{\star})$ drives this condition number to one in a neighborhood of $x_{\star}$. One can view quasi-Newton methods then as gradient descent with its preconditioner being adjusted at runtime to roughly approximate this inverse Hessian. The work [26] introduces a multidimensional backtracking technique to update a diagonal preconditioner at runtime. As a modern advancement of this idea, the Online Gradient Scaling Method of [17, 10] uses the online learning regret bound toolbox to adjust the preconditioner at runtime. Asymptotically, this ensures an optimal preconditioned performance measure for gradient descent.
@@
While effective, a downside of backtracking is that it can be “wasteful” with respect to oracle evaluations. When a backtrack occurs, the new oracle information computed at the candidate iterate is effectively discarded. Linesearch-free methods manage to avoid this overhead through various techniques. In bundle methods [24, 25, 31, 28], this is handled with a “null step” in which the function value and gradient information are still saved in the bundle to inform future iterations. Starting with the introduction of adaptive gradient descent by [33], many recent works [50, 34, 32, 30, 53, 45, 27] adaptively learn the smoothness constant without backtracking, while also providing a non-asymptotic convergence guarantee. The task of designing methods that estimate an unknown level of strong convexity is typically much harder than adapting to smoothness. Recently, [8] introduced a method for adaptively learning strong convexity constants while guaranteeing (suboptimal) linear global convergence. Our ASPGM will employ a similar estimate as a heuristic to benefit from strong convexity.
@@ 1.2 Our Contributions
  • An Adaptive Subgame Perfect Gradient Method (ASPGM). The strong non-asymptotic convergence guarantees and certificates resulting from BSPGM’s subgame perfection provide a robust restarting condition for the algorithm, ensuring a contraction under strong convexity. Combining this restarting with preconditioning gives our full algorithm, ASPGM.
@@ 2 Preliminaries
where $f$ is convex and differentiable with gradient $\nabla f$ locally Lipschitz and with a minimizer $x_{\star}$. We highlight the fact that a global Lipschitz constant $L$ on the gradient need not be known, or even exist, for our algorithm to be well-defined and progress. We denote $I$ as the identity matrix, $e_{i}$ as the $i$th standard basis vector, and $\mathbf{1}$ as the all ones vector, where the dimension of each should be clear from context.
@@
$\tilde{L}_{B}(x,y)=\frac{\frac{1}{2}\|{\nabla_{\!\!{\scriptscriptstyle B}}f(x)-\nabla_{\!\!{\scriptscriptstyle B}}f(y)}\|_{B}^{2}}{f(y)-f(x)-\langle{\nabla_{\!\!{\scriptscriptstyle B}}f(x)},{y-x}\rangle_{B}}$ (2.5)
@@ 2.1.1  OGM - An Inductive View
for some $\tau_{n}>0$ and iterate sequences $x_{n},z_{n+1}$. The base case of this induction can be established for any given $x_{0}$ by setting $z_{1}=x_{0}-\frac{2}{L}g_{0},\ \tau_{0}=2$ since we then have $H_{0}=2Q_{\star,0}(L)\geq 0.$ Then OGM iterates by setting
@@ 2.1.2  SPGM - A Dynamically Optimized Induction
$q_{i}$ $:=\tau_{i}(f_{i}-\frac{1}{2L}\|{g_{i}}\|_{B}^{2})+\frac{L}{2}\|{z_{i+1}-x_{0}}\|_{B}^{2}-v_{m}\tau_{i}$
$r_{i}$ $:=f_{i}-\langle{g_{i}},{x_{i}-x_{0}}\rangle_{B}+\frac{1}{2L}\|{g_{i}}\|_{B}^{2}-v_{m}.$
@@ 2.1.3  OBL - A Backtracking-Friendly Induction
Both of the above methods rely on knowledge of the global smoothness constant $L$, which can be both unrealistic in practice and lead to overly conservative algorithms. One classic approach to avoiding this is the use of backtracking, where at each iteration an estimate $L_{n}$ is utilized to compute a step. Then, whichever inequalities are needed by the proof are checked; if they hold the step is accepted, otherwise $L_{n}$ is increased, typically to $2L_{n}$, and the process repeats.
@@
Following this, one has that (see [46, Theorem 6])
@@ 3.1 Dynamically Optimized Induction
We begin by introducing some notation. Given $L_{n}>0$, define the vector $v\in\mathbb{R}^{k}$ by its components:
@@
for $i\in[n-k,n-1]$. Next, given a vector $\tau=(\tau_{n-k},\dots,\tau_{n-1})$ and letting $J=\{i\in[n-k,n-1]\mid\tau_{i}>0\}$, define two special indices $m$ and $s$ according to
@@
$a_{i}$ $:=\tau_{i}(f_{i}-\frac{1}{2L_{i}}\|{g_{i}}\|_{B}^{2})+\frac{L_{i}}{2}\|{z_{i+1}-x_{0}}\|_{B}^{2}-v_{m}\tau_{i}$ (3.5)
@@
Let $Z\in\mathbb{R}^{d\times k}$ be the matrix with columns given by $\frac{L_{i}}{L_{n}}(z_{i+1}-x_{0})$ and indexed by [nk,n1][n-k,n-1]. Similarly, let $G\in\mathbb{R}^{d\times k}$ be the matrix with columns now given by $\frac{g_{i}}{L_{n}}$ and indexed by [nk,n1][n-k,n-1]. Once again, by rearranging (3.8) and applying a bound on $\Delta^{\prime}$, we find that the following constraints must be satisfied:
@@
Observe that maximizing $\tau^{\prime}$ in (3.12) once again reduces to solving a low-dimensional second order cone program (SOCP). We emphasize that this SOCP is independent of the problem dimension $d$. Specifically, the subproblem (3.12) has dimension $2k$, where $k$ is the chosen memory size. By appropriately tracking and storing $O(k^{2})$ past vector-vector products (i.e., $Z^{T}Z,G^{T}G,G^{T}Z$), we eliminate any matrix-vector products with full-dimension matrices. Additionally, the problem has a linear objective with a single quadratic constraint, further reducing the complexity of the subproblem. Thus, for high-dimensional ($k\ll d$) problems with moderate costs for gradient and/or function evaluations, the per iteration computational cost of (3.12) becomes negligible.
@@ 3.2 A Backtracking-free Subgame Perfect Gradient Method (BSPGM)
This dynamic induction already suggests a reasonable backtracking method. Given $\tau^{\prime}$ and $z^{\prime}$ after solving (3.12), generate $x_{n}$ (along with $\tau_{n},z_{n+1},\Delta_{n}$) according to (2.16). We then check if the current estimate of $L_{n}$ is valid, i.e., if $Q_{m,n}(L_{n})\geq 0$. If so, accept $x_{n}$ and continue. If not, reject $x_{n}$, increase $L_{n}\leftarrow 2L_{n}$, and repeat the update rule. Note that when $x_{n}$ is accepted, this indicates that our hypothesis $U_{n}\geq 0$ is valid, so the induction can continue.
While this approach represents an effective algorithm, it is inefficient in terms of how it uses new oracle information. When a proposed next iterate $x_{n}$ is rejected, the new gradient $g_{n}$ and function value $f_{n}$ are effectively discarded without providing additional information beyond the fact that the current estimate $L_{n}$ is too small. For high-dimensional problems, first-order oracles can have a large computational cost; we therefore want to learn as much information as possible from each oracle call.
@@
This strategy is similar to the typical bundle method approach of allowing both “serious steps” where the main iterate sequence updates and “null steps” where only the bundle of first-order information is improved. Algorithm˜4 formalizes our use of this conditional approach. We emphasize that while there are remnants of a backtracking mechanism in how we iteratively increase $L_{n}$, this method is notionally distinct as we are not discarding any computed $x_{n}$ and associated first-order oracle values when $L_{n}$ increases. Rather, $(x_{n},f_{n},g_{n})$ are incorporated into our bundle of historical information via the nonnegative quantity $W_{\star,n}$. We also note the use of $\tilde{L}_{B}(x_{m},x_{n})$ to increase $L_{n}$ matches the auto-conditioning approach used by [33, 32, 50] and others.
@@ 3.3 Convergence Guarantees and Subgame Perfection of BSPGM
To maintain our inductive hypothesis $U_{n}\geq 0$, it suffices to verify that
@@
$f(x_{n})-\frac{1}{2L_{n}}\|g_{n}\|^{2}_{B}-f(x_{\star})$ $\leq\frac{\frac{L_{n}}{2}\|x_{0}-x_{\star}\|^{2}_{B}+\Delta_{n}}{\tau_{n}}\qquad\qquad$ $\text{if\ }n<N$ (3.14)
$f(x_{N})-f(x_{\star})$ $\leq\frac{\frac{L_{N}}{2}\|x_{0}-x_{\star}\|^{2}_{B}+\Delta_{N}}{\tau_{N}}\qquad\qquad$ $\text{if\ }n=N.$ (3.15)
@@
The final bound above is minimax optimal among all gradient-span methods for minimization against such an oracle and is also attained by the simpler OBL method, see [46, Theorem 4]. That is, letting $\mathcal{A}$ denote the set of gradient span methods producing a final iterate after $N$ iterations,
@@ 3.4 Adaptive Restart
$\tau_{n}\geq\frac{2{L}^{(\ell)}_{n}}{{\mu}^{(\ell)}}+\frac{2{\Delta}^{(\ell)}_{n}}{f({x}^{(\ell)}_{0})-f({x}^{(\ell)}_{n})},\qquad\qquad f({x}^{(\ell)}_{0})-f({x}^{(\ell)}_{n})>0.$ (3.16)
@@
$n\geq\sqrt{\frac{4{L}^{(\ell)}_{n}}{\mu}+\frac{4{\Delta}^{(\ell)}_{n}}{f({x}^{(\ell)}_{0})-f(x_{\star})}}+\log_{2}\left(\frac{{L}^{(\ell)}_{n}}{{L_{0}}^{(\ell)}}\right).$
@@
In practice, accurately estimating ${\mu}^{(\ell)}$ (at each epoch) can be difficult and lead algorithms to be overly conservative. Numerically, we found strong practical performance by constructing an online estimate of $\mu$ similar to the approach of [8]. We initially set ${\mu}^{(\ell)}_{0}$$\infty$ and update at each iteration according to